On July 1, 2019, Sandhill Co. purchased new equipment for $90,000. Its estimated useful life was 8 years with a $18,000 salvage value. On January 1, 2022, before making its depreciation entry for 2022, the company estimated the remaining useful life to be 10 years beyond December 31, 2022. The new salvage value is estimated to be $5,000.

a.Prepare the journal entry to record depreciation on December 31, 2019. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)
b. Prepare the journal entry to record depreciation on December 31, 2020. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

c. Compute the revised annual depreciation on December 31, 2022.
d. Prepare the journal entry to record depreciation on December 31, 2022. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)
e. Compute the balance in Accumulated Depreciation—Equipment for this equipment after depreciation expense has been recorded on December 31, 2022.
Answer #1

a) Journal entry

Date account and explanation Debit credit
Depreciation expense (90000-18000/8)*6/12 4500
Accumulated depreciation-equipment 4500

b) Journal entry

Date account and explanation Debit credit
Depreciation expense (90000-18000/8) 9000
Accumulated depreciation-equipment 9000

c) Revised annual depreciation = (90000-22500-5000)/10 = 6250

d) Journal entry

Date account and explanation Debit credit
Depreciation expense 6250
Accumulated depreciation-equipment 6250

e) Accumulated depreciation-equipment = 22500+6250 = 28750